OpenStudy (anonymous):

find the remainder when p(x) = x3 + 3x2 - 2x + 4 is divided by 2x + 1

OpenStudy (anonymous):

if u want to find it without division, put x=-1/2 in the polynomial

OpenStudy (anonymous):

(-1/2)^3+3(-1/2)^2-2(-1/2)+4=-1/8+3/4+3=-1+6+24/8=29/8

OpenStudy (dumbcow):

use long division .5x^2 + 1.25x - 1.625 2x+1 |x^3 +3x^2 -2x +4 -( x^3 +.5x^2) _______________ 2.5x^2 -2x -(2.5x^2 +1.25x) _________________ -3.25x +4 -(-3.25x -1.625) _________________ 5.625 remainder is 5.625